For each of the following numbers, find the smallest whole number by which it should be divided so as to get a perfect square. Also find the square root of the square number so obtained. i. 252 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

252 = 2×2×3×3×7
= (2×2)×(3×3)×7
Here, 7 cannot be paired.
∴ We will divide 252 by 7 to get perfect square. New number = 252 ÷ 7= 36
square root of 36= 6
Thanks